A particle moves along the x-axis so that at time t >= 0 (t is greater than or equal to "t") its position is given by x(t)=-t^3+11t^2-24t. Determine the velocity of the particle at t=7.


1
Expert's answer
2021-02-02T04:45:18-0500

"x(t) = -t^3 +11t^2 -24t \\

v(t) = xʼ(t) = -3t^2+22t -24"

At t = 7

"v(7) = -3 \times 7^2 + 22 \times 7 -24 \\

= -147 +154 -24 \\

= -17"

Answer: -17